5.14 Find the future values of these ordinary annuities. Compounding occurs once a year. Do not...

5.14

Find the future values of these ordinary annuities. Compounding occurs once a year. Do not round intermediate calculations. Round your answers to the nearest cent.

  1. $300 per year for 16 years at 12%.

    $  

  2. $150 per year for 8 years at 6%.

    $  

  3. $300 per year for 8 years at 0%.

    $  

  4. Rework parts a, b, and c assuming they are annuities due.

    Future value of $300 per year for 16 years at 12%: $  

    Future value of $150 per year for 8 years at 6%: $  

    Future value of $300 per year for 8 years at 0%: $  
Answer #1

Future value of annuity=Annuity[(1+rate)^time period-1]/rate

a.Future value=300[(1.12)^16-1]/0.12

=300*42.75328042

=12825.98(Approx).

b.Future value=150[(1.06)^8-1]/0.06

=150*9.897467909

=$1484.62(Approx).

c.Future value=300*8

=2400

Future value of annuity due=Future value of annuity*(1+rate)

a.Future value=12825.98*1.12

=14365.10(Approx).

b.Future value=1484.62*1.06

=$1573.70(Approx).

c.Future value=300*8

=2400